OpenStudy (lovelyharmonics):

Find the zeros of the polynomial function and state the multiplicity of each. f(x) = 3(x + 8)2(x - 8)3

OpenStudy (lovelyharmonics):

@surjithayer ._. im sorry to keep bothering you

OpenStudy (anonymous):

Refer to the attached plot.

OpenStudy (lovelyharmonics):

what is that? o.O

OpenStudy (anonymous):

put x+8=0,x=-8 isa zero of multiplicity 2 x-8=0,x=8 is a zero of multiplicity 3

OpenStudy (lovelyharmonics):

okay but what about the 3 out in front?

OpenStudy (lovelyharmonics):

if you could, i have a few other questions that i need help with too c:

OpenStudy (anonymous):

The attachement is a plot of\[x^5-8 x^4-128 x^3+1024 x^2+4096 x-32768 \]

OpenStudy (lovelyharmonics):

Find a cubic function with the given zeros. Square root of seven., -Square root of seven., -4 how are you so good at math?(◕︵◕)

OpenStudy (anonymous):

\[f(x)=\left( x-\sqrt{7} \right)\left( x+\sqrt{7} \right)\left( x+4 \right)=?\] solve first two first

OpenStudy (lovelyharmonics):

theyd be the opposite so -sqrt7 would be sqrt7 and sqrt7 would be -sqrt7

OpenStudy (anonymous):

\[x=\sqrt{7} ~is`~a ~zero~of~f(x)~so~x-\sqrt{7}~is ~a~factor~of~f(x)\]

OpenStudy (lovelyharmonics):

wait what? 0.o

OpenStudy (anonymous):

This is the reason for changing sign

OpenStudy (lovelyharmonics):

changing the sign? didnt you just subtract sqrt7 from both sides?

OpenStudy (anonymous):

\[\left( a+b \right)\left( a-b \right)=a^2-b^2\] x^2+2x-3=(x+3)(x-1) it has zeros -3 and 1 but x+3 and x-1 are factors

OpenStudy (anonymous):

\[\left( x+\sqrt{7} \right)\left( x-\sqrt{7} \right)=x^2-(\sqrt{7})^2=x^2-7\]

OpenStudy (lovelyharmonics):

okay, but i was right? so -sqrt7 and sqrt7 are 2 of the answers

OpenStudy (anonymous):

correct ,they are zeros.

OpenStudy (lovelyharmonics):

yay c: okay so what about the tricky last part?

OpenStudy (lovelyharmonics):

wait isnt it just -4?

OpenStudy (anonymous):

solve\[\left( x^2-7 \right)\left( x+4 \right)\]

OpenStudy (anonymous):

solve means multiply.

OpenStudy (lovelyharmonics):

x^3+4x^2-7x-28

OpenStudy (anonymous):

correct

OpenStudy (lovelyharmonics):

okay one more question c: State the domain of the rational function. f(x) = seven divided by quantity fourteen minus x.

OpenStudy (anonymous):

\[f(x)=\frac{ 7 }{14-x }\] it is defined for all values except where denominator is 0 for domain\[14-x \neq 0~ or~x \neq14\] domain is all real values except x=14

OpenStudy (lovelyharmonics):

its that easy? o.o

OpenStudy (anonymous):

yes

OpenStudy (lovelyharmonics):

okay thank you c:

OpenStudy (anonymous):

yw